Mrs. Kwok is 6 years more than 3 times as old as her daughter. 8 years ago she was 2 years less than 9 times as old as her daughter was then. Find each of their ages.

The sum of 3 times John's age and 2 times jill's age is 44. Jill is 8 years less than twice as old as John is. Find their ages

Answer by Photonjohn(42)

D = age of daughter
K = age of mother
3D + 6 = K
8 years ago,
K-8+2 = 9(D-8)
3D+6-8+2 = 9D-72
6D=72
D=12
K=12x3+6 = 42
J = John's age
L = Jill's age
3J + 2L = 44
L=8+2J
J=4
L=16